mp/1799965-154 704032381x+ plane Answer in units of m/s2. 019 10.0 points A bee flies to a flower 569 m due south of its hive. The bees speed in still air is 0.73 m/s, and there is a wind blowing toward the south at 0.15 m/s. How long will it take the bee to travel to the flower and back to the hive? Answer in units of s
Answer #1

Bee's speed in still air Vb = 0.73 m/s

Wind blowing towards south with a speed, Vw = 0.15 m/s

So, when the bee travels towards the flower, its speed = Vb + Vw = 0.73 + 0.15 = 0.88 m/s

So, time taken in the upward journey, t1 = distance/speed = 569 / 0.88 = 646.6 s

Now, for downward journey back to home -

its speed = Vw - Vb = 0.73 - 0.15 = 0.58 m/s

Time taken in the downward journey, t2 = 569 / 0.58 = 981.0 s

Therefore, total time of journey = t1 + t2 = 646.6 + 981.0 = 1627.6 s